OpenStudy (anonymous):

If f(x) = 2x2 + 2, find f(5).

OpenStudy (nubeer):

Subsitute x with 5 and solve it

OpenStudy (anonymous):

\[f(x) = 2x^2 + 2 \rightarrow f(5) = 2(5)^2 + 2 =2 \times 25 +2 = 50+2 =52\] Thus f(5) = 52